# A Survey of Deep Learning for Geometry Problem Solving

### v3

@@ 1 Introduction
Figure 1: Overview of the survey‘s structure
@@
In this study, we started with classic papers in this field, used a single round of forward and backward snowballing, searched Google Scholar with the keyword "geometry", and manually screened to ensure the relevance of the papers. In the end, we collected more than 310 academic papers that involved deep learning for geometry problem solving, and conducted a comprehensive and in-depth survey.
@@ 4.1.1 Performance-Based Metrics
Outcome-Based Metrics. Outcome-based metrics focus on measuring the correctness of the final inference results without considering the details of the reasoning process. Top-k Accuracy (Top-k Acc) and Pass@n (P@n) are two main metrics for measuring answer accuracy. Top-k Acc refers to the ratio of the first k predictions generated by the model that have at least one correct answer, while P@n refers to the ratio of the model that can correctly solve the problem at least once in a given n attempts. Other works also use outcome-based metrics such as F1 score Mishra et al. (2022b); Cheng et al. (2025b), maj@k Yue et al. (2024a), number of correct and wrong answers Dou et al. (2024), choice Zhang et al. (2023a) and competition scores. F1 score is a metric that takes into account both precision and recall; maj@k measures the probability that the model gets the correct answer through majority vote in k samples; the number of correct and wrong answers intuitively shows the task results; choice refers to the probability of selecting the correct answer from the candidate options of a multiple-choice question and if it cannot be determined, an option is selected randomly; some studies also show model performance by showing SAT scores Seo et al. (2015) or IMO competition scores Trinh et al. (2024).
@@ 4.1.2 Efficiency-Based Metrics
Efficiency-based metrics are used to measure the resource consumption and efficiency performance of the model during reasoning. These metrics include the time it takes to solve the problem Alvin et al. (2017), the ratio of unsuccessful solutions within the specified time (timeout) Zhang et al. (2023c), the number of inference steps Wu et al. (2024a); Fang et al. (2024), and the cost of running the model Balunović et al. (2025).
@@ 4.2.3 LLM-as-a-Judge
Recently, as the capabilities of LLMs continue to increase, more and more studies have begun to use the "LLM-as-a-judge" paradigm for evaluation rather than rule-based methods Li et al. (2024a). This paradigm is not only more cost-effective than manual evaluation, but also more scalable than rule-based evaluation, and is suitable for complex tasks such as geometry problem solving. Common uses include extracting answers Lu et al. (2024); Zhang et al. (2024e), standardizing answers Gao et al. (2025a), evaluating the reasoning process Anand et al. (2024a); Jaiswal et al. (2024), and checking errors Hendrycks et al. (2021); Sun et al. (2024). When evaluating the quality of the reasoning process, because some questions may have multiple solutions, rigid standard answers may limit the accuracy of actual reasoning evaluation. Using LLM to extract the key steps of the reasoning process and score them one by one is more flexible and more in line with real scenarios Zhang et al. (2024g). However, we need to be wary of the possible hallucinations and extra resource consumption that LLM may cause.
@@ 5.1 Challenges
Evaluation. First, the question types are monotonous. Existing benchmarks mainly use multiple-choice questions for evaluation (see Table 1), which provides space for models to guess answers, thus affecting the accuracy and reliability of the evaluation results. Although some works have alleviated this problem by evaluating different permutations of options to reduce the chance error (CircularEval) Liu et al. (2024b) or by not providing candidate options Fu et al. (2025), these methods have not yet been widely used. Second, there is no standard method for evaluating the reasoning process. Recently, with the increasing demand for model improvement, the evaluation of the geometry reasoning process has gradually received attention. However, these evaluation methods lack unified standards, and we should formulate more precise evaluation standards to accurately identify and improve the deficiencies of the model Park et al. (2024). In addition, current benchmarks may lack robustness, and the performance of the model is often not stable enough after applying small perturbations to the same evaluation questions Wang et al. (2025d); Zhou et al. (2024d); Weng et al. (2024). At the same time, some datasets may be included in the training data, thus affecting the true evaluation of model performance Park et al. (2024). This requires the academic community to introduce a fairer and more authoritative evaluation method.
@@ 5.2 Future Directions
Using Cognitive Pattern. Cognitive pattern is a comprehensive approach that aims to simulate the cognitive process of humans in understanding and solving complex problems Kurbatov et al. (2021, 2022). The concept appeared in early problem-solving research. In the study of geometry problem solving, many strategies that simulate the human problem-solving process have been proven to be effective Zhou and Yu (2021); Rao et al. (2022), including: highlighting key information in diagrams and texts to focus on the key points of the problem; referencing diagram annotations, or adding auxiliary lines, coordinate axes and other elements to the diagram to enhance the understanding of the geometric structure; searching and applying relevant geometry theorems and knowledge to guide the problem-solving process; using curriculum learning methods to improve the model’s problem-solving ability gradually. These methods have not yet been fully integrated into current advanced deep learning systems and deserve further attention and exploration.

### v4

@@ 1 Introduction
Figure 1: Overview of the survey’s structure
@@
In this study, we started with classic papers in this field, used a single round of forward and backward snowballing, searched Google Scholar with the keyword “geometry”, and manually screened to ensure the relevance of the papers. In the end, we collected more than 310 academic papers that involved deep learning for geometry problem solving, and conducted a comprehensive and in-depth survey.
@@ 4.1.1 Performance-Based Metrics
Outcome-Based Metrics. Outcome-based metrics focus on measuring the correctness of the final inference results without considering the details of the reasoning process. Top-k Accuracy (Top-k Acc) and Pass@n (P@n) are two main metrics for measuring answer accuracy. Top-k Acc refers to the ratio of the first k predictions generated by the model that have at least one correct answer, while P@n refers to the ratio of the model that can correctly solve the problem at least once in a given n attempts. Other works also use outcome-based metrics such as F1 score Mishra et al. (2022b); Cheng et al. (2025b), maj@k Yue et al. (2024a), number of correct and wrong answers Dou et al. (2024), choice Zhang et al. (2023a) and competition scores. F1 score is a metric that takes into account both precision and recall; maj@k measures the ratio of cases that the model gets the correct answer through majority vote in k samples; the number of correct and wrong answers intuitively shows the task results; choice refers to the ratio of correctly selected answers from four candidate options of a multiple-choice question, with a random selection made if the model’s output does not match any of the given options; some studies also show model performance by showing SAT scores Seo et al. (2015) or IMO competition scores Trinh et al. (2024).
@@ 4.1.2 Efficiency-Based Metrics
Efficiency-based metrics are used to measure the resource consumption and efficiency performance of the model during reasoning. These metrics include the time it takes to solve the problem Alvin et al. (2017), the ratio of unsuccessful solutions within the specified time (timeoutZhang et al. (2023c), the number of inference steps Wu et al. (2024a); Fang et al. (2024), and the cost of running the model Balunović et al. (2025).
@@ 4.2.3 LLM-as-a-Judge
Recently, as the capabilities of LLMs continue to increase, more and more studies have begun to use the “LLM-as-a-judge” paradigm for evaluation rather than rule-based methods Li et al. (2024a). This paradigm is not only more cost-effective than manual evaluation, but also more scalable than rule-based evaluation, and is suitable for complex tasks such as geometry problem solving. Common uses include extracting answers Lu et al. (2024); Zhang et al. (2024e), standardizing answers Gao et al. (2025a), evaluating the reasoning process Anand et al. (2024a); Jaiswal et al. (2024), and checking errors Hendrycks et al. (2021); Sun et al. (2024). When evaluating the quality of the reasoning process, because some questions may have multiple solutions, rigid standard answers may limit the accuracy of actual reasoning evaluation. Using LLM to extract the key steps of the reasoning process and score them one by one is more flexible and more in line with real scenarios Zhang et al. (2024g). However, we need to be wary of the possible hallucinations and extra resource consumption that LLM may cause.
@@ 5.1 Challenges
Evaluation. First, the question types are monotonous. Existing benchmarks mainly use multiple-choice questions for evaluation (see Table 1), which provides space for models to guess answers, thus affecting the accuracy and reliability of the evaluation results. Although some studies have attempted to alleviate this issue—such as by evaluating different permutations of answer choices to mitigate positional bias (as in CircularEval) Liu et al. (2024b), or by removing candidate options entirely Fu et al. (2025)—these strategies have not yet been widely adopted. Second, there is no standard method for evaluating the reasoning process. Recently, with the increasing demand for model improvement, the evaluation of the geometry reasoning process has gradually received attention. However, these evaluation methods lack unified standards, and we should formulate more precise evaluation standards to accurately identify and improve the deficiencies of the model Park et al. (2024). In addition, current benchmarks may lack robustness, and the performance of the model is often not stable enough after applying small perturbations to the same evaluation questions Wang et al. (2025d); Zhou et al. (2024d); Weng et al. (2024). At the same time, some datasets may be included in the training data, thus affecting the true evaluation of model performance Park et al. (2024). This requires the academic community to introduce a fairer and more authoritative evaluation method.
@@ 5.2 Future Directions
Using Cognitive Pattern. Cognitive pattern is a comprehensive approach that aims to simulate the cognitive process of humans in understanding and solving complex problems Kurbatov et al. (2021, 2022). The concept appeared in early problem-solving research. In the study of geometry problem solving, many strategies that simulate the human problem-solving process have been proven to be effective Zhou and Yu (2021); Rao et al. (2022). These strategies include: highlighting key information in diagrams and texts to focus on the key points of the problem; referencing diagram annotations, or adding auxiliary lines, coordinate axes and other elements to the diagram to enhance the understanding of the geometric structure; searching and applying relevant geometry theorems and knowledge to guide the problem-solving process; and using curriculum learning methods to improve the model’s problem-solving ability gradually. These methods have not yet been fully integrated into current advanced deep learning systems and deserve further attention and exploration.